OpenStudy (anonymous):

What is the best way to classify each equation? 6(2x-3)=3(3x-5) A. identity B. contradiction C. neither

OpenStudy (anonymous):

@chosenmatt @shinalcantara

OpenStudy (shinalcantara):

6(2x - 3) =3(3x -5) 12x - 18 = 9x - 15 12x - 9x = -15 + 18 3x = 3 x = 1 there's a solution but the left and right side of the equation are not equal . it's C